OpenStudy (anonymous):

A single household circuit is connected to three electrical outlets (all in parallel). The outlets are connected to three devices: an 800-W electric kettle, a 60-W trouble light, and a 1,000-W toaster oven. If all three appliances operate at the same time, will the 15-amp fuse of the circuit trip? yes or no ?

OpenStudy (radar):

First calculate the total wattage. Addition

OpenStudy (radar):

Next make an assumption: That all devices are 120 volt units, and the outlets furnish 120 volts. Note that this is a good assumption in the U.S.A., but would not be good in Germany or many other countries.

OpenStudy (radar):

Now dig up the formula for power, the one that relates power to volts and current (amps) P=VI or IV Now solve for I (current) I=P/V

OpenStudy (radar):

Does I exceed 15 amps (yes or no)

OpenStudy (anonymous):

Yes ?

OpenStudy (radar):

Total watts being consumed: 800 + 60 +1000 = 1860 watts I=1860/120 = 15.5 Amperes that exceeds the rating of 15, so yes fuse opens up or circuit breaker will trip open.

OpenStudy (anonymous):

thank you sir

OpenStudy (radar):

You're welcome, and good luck with your studies.